what is the equation for the circumscribed circle whose triangle verticies are A = (-5,12), B = (5, -12), and C = (5,12)?
thanks

Answer by scott8148(6628)

A and C have the same y coordinate (12), so AC is a vertical line

B and C have the same x coordinate (5), so BC is a horizontal line

AC and BC are perpendicular, so ABC is a right triangle
__ this means that the hypotenuse (AB) is the diameter of the circumscribed circle

the center of the circle is the midpoint of AB __ (5-5)/2 and (12-12)/2 __ (0,0) or the origin

the radius of the circle is the distance from the center to either A or B
__ by Pythagoras, r^2=(-5)^2+12^2 __ r^2=169 __ r=13

so the equation of the circle is (x-0)^2 + (y-0)^2 = 13^2 __ x^2+y^2=169